ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

19. History of Mathematics

1991 MS Classification: 1 Coonce, Harry B., Mankato State University, Mankato, MN USA The Genealogy Project for Mathematicians The Genealogy Project is an attempt to trace the development of mathematics via the advisor/advisee relationship. As of April 1, 1998 over 10,000 such connections have been identified. When ideal conditions are met we show for each individual the following: Name, university where the degree was earned, year of the degree, the dissertation title and the name of the advisor(s). We present the results of this work as we obtain and process the data on the World Wide Web. Links are provided to the advisor of each person and to each of his students. Information will be given regarding how one can participate. This project is joint work with Steve Butcher and Pat Allaire. Reference: http://hcoonce.math.mankato.msus.edu